# xRFM: Accurate, scalable, and interpretable feature learning models for tabular data

### v1

@@ Appendix B All results
RealMLP 7.26 0.839 0.952 7.41 14.8 33.3 74.1 0.125
TabR 7.72 0.844 0.971 18.5 33.3 55.6 70.4 0.116
xRFM 8.83 0.841 0.958 14.8 25.9 37.0 66.7 0.123
ModernNCA 9.50 0.843 0.965 3.70 33.3 44.4 59.3 0.120
CatBoost 9.93 0.827 0.892 11.1 29.6 37.0 55.6 0.128
LightGBM 10.2 0.826 0.886 7.41 25.9 44.4 55.6 0.128
TabPFN-v2 10.8 0.834 0.924 14.8 22.2 29.6 48.1 0.129
MLP-PLR 11.2 0.827 0.891 0.0 3.70 18.5 40.7 0.131
XGBoost 11.9 0.824 0.876 7.41 22.2 33.3 51.9 0.129
FTT 12.3 0.829 0.902 0.0 11.1 18.5 33.3 0.135
MLP 12.6 0.833 0.916 0.0 7.41 14.8 29.6 0.130
ResNet 13.1 0.834 0.921 0.0 7.41 7.41 29.6 0.130
DCN2 13.2 0.832 0.920 0.0 3.70 3.70 25.9 0.130
PTARL 13.3 0.831 0.912 0.0 3.70 11.1 14.8 0.131
AutoInt 13.6 0.831 0.912 0.0 0.0 3.70 14.8 0.130
Excelformer 14.6 0.821 0.859 3.70 7.41 7.41 11.1 0.143
SNN 14.6 0.831 0.913 0.0 3.70 7.41 18.5 0.131
Laplace KRR 17.0 0.826 0.881 3.70 11.1 14.8 18.5 0.138
Danets 17.3 0.827 0.889 0.0 0.0 0.0 0.0 0.133
TabCaps 18.4 0.822 0.863 0.0 3.70 7.41 7.41 0.136
TabTransformer 19.2 0.816 0.797 0.0 3.70 7.41 18.5 0.160
RandomForest 19.5 0.806 0.773 0.0 3.70 7.41 18.5 0.144
Node 20.1 0.793 0.725 3.70 3.70 3.70 7.41 0.151
LogReg 20.5 0.807 0.779 0.0 7.41 18.5 25.9 0.159
Tangos 20.7 0.811 0.799 0.0 0.0 3.70 3.70 0.139
SVM 21.8 0.802 0.755 0.0 0.0 3.70 18.5 0.162
GrowNet 22.6 0.804 0.772 0.0 0.0 0.0 3.70 0.160
TabNet 23.2 0.814 0.819 0.0 0.0 0.0 0.0 0.142
kNN 24.3 0.787 0.674 0.0 0.0 0.0 7.41 0.166
SwitchTab 27.3 0.786 0.671 0.0 0.0 0.0 0.0 0.163
@@
TabPFN-v2 4.99 0.864 0.946 25.8 62.4 74.2 80.6 0.104
ModernNCA 10.2 0.856 0.886 14.0 21.5 33.3 51.6 0.111
CatBoost 10.2 0.845 0.867 5.38 23.7 35.5 55.9 0.119
LightGBM 10.3 0.845 0.869 6.45 19.4 35.5 55.9 0.119
TabR 10.4 0.851 0.872 4.30 14.0 30.1 44.1 0.116
xRFM 10.7 0.853 0.869 5.38 29.0 41.9 51.6 0.114
XGBoost 11.3 0.844 0.850 4.30 17.2 34.4 49.5 0.120
RealMLP 11.5 0.850 0.863 1.08 7.53 18.3 37.6 0.116
FTT 13.2 0.836 0.813 1.08 7.53 18.3 33.3 0.124
MLP-PLR 14.1 0.838 0.816 0.0 3.23 11.8 26.9 0.124
RandomForest 14.2 0.836 0.823 4.30 10.8 20.4 34.4 0.129
AutoInt 15.1 0.834 0.803 0.0 3.23 5.38 16.1 0.129
Tangos 15.3 0.834 0.799 0.0 2.15 7.53 18.3 0.131
DCN2 15.5 0.839 0.815 1.08 2.15 9.68 21.5 0.127
Laplace KRR 15.6 0.836 0.800 2.15 15.1 20.4 26.9 0.133
MLP 15.9 0.836 0.796 1.08 3.23 8.60 17.2 0.132
ResNet 16.0 0.839 0.810 1.08 3.23 8.60 17.2 0.129
TabCaps 16.1 0.835 0.804 1.08 3.23 6.45 14.0 0.130
Excelformer 16.5 0.830 0.778 0.0 1.08 7.53 19.4 0.131
SNN 16.6 0.836 0.800 1.08 3.23 4.30 6.45 0.129
Node 16.9 0.824 0.756 0.0 6.45 15.1 25.8 0.137
PTARL 16.9 0.830 0.779 0.0 1.08 3.23 7.53 0.134
Danets 17.9 0.829 0.758 0.0 3.23 7.53 15.1 0.137
TabTransformer 19.7 0.820 0.714 1.08 1.08 7.53 12.9 0.148
LogReg 20.1 0.819 0.723 3.23 7.53 12.9 18.3 0.146
kNN 20.8 0.811 0.677 5.38 9.68 14.0 18.3 0.158
SVM 21.5 0.816 0.697 2.15 4.30 6.45 11.8 0.150
GrowNet 22.1 0.817 0.713 0.0 0.0 0.0 2.15 0.151
SwitchTab 23.9 0.809 0.675 0.0 0.0 0.0 1.08 0.156
TabNet 25.8 0.802 0.633 0.0 0.0 0.0 0.0 0.157
NaiveBayes 28.6 0.695 0.280 1.08 3.23 3.23 4.30 0.253
NCM 29.9 0.723 0.231 0.0 0.0 1.08 2.15 0.257
@@
Dataset xRFM XGB CatBoost LGBM RealMLP MLP-PLR MLP ResNet
Airlines_DepDelay_10M 0.9823 0.9813 0.9796 0.9798 0.9786 0.9795 0.9824 0.9818
Allstate_Claims_Severity 0.6544 0.6547 0.6510 0.6530 0.6495 0.6537 0.6557 0.6537
black_friday 0.6864 0.6807 0.6792 0.6787 0.6859 0.6862 0.6929 0.6892
Buzzinsocialmedia_Twitter 0.2035 0.2134 0.3147 0.2789 0.2566 0.2553 0.2840 0.2906
nyc-taxi-green-dec-2016 0.5817 0.6649 0.6567 0.6489 0.6142 0.6523 0.6657 0.6365
wave_energy 0.0022 0.0918 0.0499 0.0821 0.0024 0.0073 0.0254 0.0434
Yolanda 0.7829 0.8012 0.8094 0.7970 0.7869 0.7897 0.7927 0.7856
@@
Dataset xRFM XGB CatBoost LGBM RealMLP MLP-PLR MLP ResNet
airlines 0.3347 0.3292 0.3315 0.3287 0.3344 0.3342 0.3342 0.3339
covertype 0.0259 0.0420 0.0612 0.0333 0.0280 0.0364 0.0404 0.0385
Higgs 0.2674 0.2576 0.2576 0.2549 0.2473 0.2528 0.2515 0.2435
jannis 0.2729 0.2799 0.2816 0.2779 0.2702 0.2764 0.2865 0.2795
MiniBooNE 0.0550 0.0529 0.0538 0.0525 0.0484 0.0504 0.0503 0.0488
numerai28.6 0.4789 0.4812 0.4790 0.4782 0.4800 0.4775 0.4771 0.4800
@@
dionis 0.0945 0.1219 0.1041 0.1076 0.0887 0.1257 0.1110 0.0907
Fashion-MNIST 0.0904 0.0928 0.0969 0.0895 0.0913 0.1064 0.1041 0.1011
kick 0.0977 0.0965 0.0956 0.0964 0.0976 0.0978 0.0979 0.0970
@@
Figure B.1: Runtime comparison as a function of the number of training samples on the covertype dataset.

### v2

@@ Appendix B All results
xRFM 5.96 0.845 0.981 29.6 55.6 59.3 74.1 0.119
RealMLP 7.44 0.839 0.951 7.41 14.8 29.6 77.8 0.125
TabR 7.83 0.844 0.970 18.5 33.3 55.6 70.4 0.116
ModernNCA 9.69 0.843 0.963 3.70 29.6 40.7 59.3 0.120
CatBoost 10.2 0.827 0.891 7.41 25.9 33.3 55.6 0.128
LightGBM 10.5 0.826 0.884 7.41 18.5 40.7 55.6 0.128
TabPFN-v2 10.9 0.834 0.923 7.41 22.2 25.9 48.1 0.129
MLP-PLR 11.2 0.827 0.890 0.0 0.0 18.5 40.7 0.131
XGBoost 12.1 0.824 0.875 3.70 18.5 33.3 51.9 0.129
FTT 12.4 0.829 0.900 0.0 11.1 18.5 29.6 0.135
MLP 12.7 0.833 0.915 0.0 7.41 14.8 29.6 0.130
ResNet 13.3 0.834 0.920 0.0 3.70 7.41 29.6 0.130
DCN2 13.4 0.832 0.919 0.0 3.70 3.70 22.2 0.130
PTARL 13.4 0.831 0.911 0.0 3.70 7.41 14.8 0.131
AutoInt 13.7 0.831 0.911 0.0 0.0 3.70 11.1 0.130
Excelformer 14.7 0.821 0.858 3.70 7.41 7.41 11.1 0.143
SNN 14.8 0.831 0.912 0.0 3.70 7.41 18.5 0.131
Laplace KRR 17.0 0.826 0.879 3.70 11.1 14.8 14.8 0.138
Danets 17.4 0.827 0.887 0.0 0.0 0.0 0.0 0.133
TabCaps 18.4 0.822 0.862 0.0 3.70 7.41 7.41 0.136
TabTransformer 19.2 0.816 0.796 0.0 3.70 7.41 18.5 0.160
RandomForest 19.6 0.806 0.772 0.0 3.70 7.41 18.5 0.144
Node 20.1 0.793 0.724 3.70 3.70 3.70 7.41 0.151
LogReg 20.5 0.807 0.778 0.0 3.70 18.5 25.9 0.159
Tangos 20.7 0.811 0.798 0.0 0.0 3.70 3.70 0.139
SVM 21.8 0.802 0.754 0.0 0.0 3.70 18.5 0.162
GrowNet 22.7 0.804 0.771 0.0 0.0 0.0 3.70 0.160
TabNet 23.2 0.814 0.818 0.0 0.0 0.0 0.0 0.142
kNN 24.3 0.787 0.673 0.0 0.0 0.0 7.41 0.166
SwitchTab 27.3 0.786 0.670 0.0 0.0 0.0 0.0 0.163
@@
TabPFN-v2 5.02 0.864 0.947 26.9 59.1 72.0 80.6 0.104
xRFM 8.75 0.856 0.893 7.53 29.0 49.5 63.4 0.111
ModernNCA 10.3 0.856 0.886 11.8 22.6 31.2 50.5 0.111
CatBoost 10.3 0.845 0.867 4.30 23.7 34.4 54.8 0.119
LightGBM 10.4 0.845 0.869 7.53 18.3 35.5 54.8 0.119
TabR 10.5 0.851 0.873 4.30 17.2 30.1 43.0 0.116
XGBoost 11.4 0.844 0.851 5.38 18.3 35.5 49.5 0.120
RealMLP 11.7 0.850 0.864 1.08 7.53 18.3 35.5 0.116
FTT 13.3 0.836 0.813 1.08 6.45 17.2 32.3 0.124
MLP-PLR 14.2 0.838 0.817 0.0 3.23 11.8 29.0 0.124
RandomForest 14.3 0.836 0.824 3.23 10.8 19.4 33.3 0.129
AutoInt 15.2 0.834 0.803 0.0 3.23 5.38 16.1 0.129
Tangos 15.4 0.834 0.800 0.0 3.23 7.53 16.1 0.131
DCN2 15.5 0.839 0.816 1.08 2.15 9.68 21.5 0.127
Laplace KRR 15.7 0.836 0.801 1.08 16.1 20.4 26.9 0.133
MLP 16.0 0.836 0.797 1.08 3.23 8.60 17.2 0.132
ResNet 16.1 0.839 0.811 1.08 3.23 8.60 19.4 0.129
TabCaps 16.2 0.835 0.804 1.08 3.23 5.38 14.0 0.130
Excelformer 16.6 0.830 0.778 0.0 1.08 7.53 16.1 0.131
SNN 16.6 0.836 0.800 1.08 3.23 4.30 7.53 0.129
Node 17.0 0.824 0.757 0.0 5.38 15.1 25.8 0.137
PTARL 17.0 0.830 0.780 0.0 1.08 3.23 8.60 0.134
Danets 18.0 0.829 0.759 0.0 3.23 6.45 14.0 0.137
TabTransformer 19.7 0.820 0.715 1.08 1.08 6.45 12.9 0.148
LogReg 20.1 0.819 0.723 3.23 9.68 12.9 17.2 0.146
kNN 20.8 0.811 0.677 5.38 10.8 14.0 18.3 0.158
SVM 21.6 0.816 0.698 2.15 3.23 6.45 10.8 0.150
GrowNet 22.2 0.817 0.714 0.0 0.0 0.0 2.15 0.151
SwitchTab 24.0 0.809 0.675 0.0 0.0 0.0 0.0 0.156
TabNet 25.8 0.802 0.633 0.0 0.0 0.0 1.08 0.157
NaiveBayes 28.6 0.695 0.281 1.08 2.15 3.23 4.30 0.253
NCM 29.9 0.723 0.230 0.0 0.0 0.0 2.15 0.257
@@
Dataset xRFM XGB CatBoost LGBM RealMLP MLP-PLR MLP-RTDL ResNet-RTDL
Airlines_DepDelay_10M 0.9818 0.9813 0.9796 0.9798 0.9786 0.9795 0.9824 0.9818
Allstate_Claims_Severity 0.6489 0.6547 0.6510 0.6530 0.6495 0.6537 0.6557 0.6537
black_friday 0.6881 0.6807 0.6792 0.6787 0.6859 0.6862 0.6929 0.6892
Buzzinsocialmedia_Twitter 0.2080 0.2134 0.3147 0.2789 0.2566 0.2553 0.2840 0.2906
nyc-taxi-green-dec-2016 0.5834 0.6649 0.6567 0.6489 0.6142 0.6523 0.6657 0.6365
wave_energy 0.0020 0.0918 0.0499 0.0821 0.0024 0.0073 0.0254 0.0434
Yolanda 0.7816 0.8012 0.8094 0.7970 0.7869 0.7897 0.7927 0.7856
@@
Dataset xRFM XGB CatBoost LGBM RealMLP MLP-PLR MLP-RTDL ResNet-RTDL
airlines 0.3341 0.3292 0.3315 0.3287 0.3344 0.3342 0.3342 0.3339
covertype 0.0257 0.0420 0.0612 0.0333 0.0280 0.0364 0.0404 0.0385
Higgs 0.2640 0.2576 0.2576 0.2549 0.2473 0.2528 0.2515 0.2435
jannis 0.2701 0.2799 0.2816 0.2779 0.2702 0.2764 0.2865 0.2795
MiniBooNE 0.0539 0.0529 0.0538 0.0525 0.0484 0.0504 0.0503 0.0488
numerai28.6 0.4778 0.4812 0.4790 0.4782 0.4800 0.4775 0.4771 0.4800
@@
dionis 0.0926 0.1219 0.1041 0.1076 0.0887 0.1257 0.1110 0.0907
Fashion-MNIST 0.0889 0.0928 0.0969 0.0895 0.0913 0.1064 0.1041 0.1011
kick 0.0972 0.0965 0.0956 0.0964 0.0976 0.0978 0.0979 0.0970
@@
Model Elo ($\uparrow$) Norm. Avg. Harm. #wins ($\uparrow$) Improva- Train time Predict time
score ($\uparrow$) rank ($\downarrow$) mean bility ($\downarrow$) per 1K [s] per 1K [s]
rank ($\downarrow$)
AutoGluon 1.3 (4h) 1779 0.673 5.2 3.2 1 3.6% 1734.20 7.06
RealMLP (T+E) 1721 0.660 6.6 5.4 0 3.1% 6860.54 7.68
ModernNCA (T+E) 1626 0.622 9.2 2.6 3 5.4% 3811.43 7.58
LightGBM (T+E) 1602 0.478 10.1 7.5 0 6.0% 686.46 5.48
TabDPT (D) 1575 0.515 10.9 3.8 2 3.9% 16.97 8.70
xRFM (T+E) 1563 0.492 11.5 5.3 1 5.1% 365.57 0.72
CatBoost (T+E) 1558 0.435 11.8 8.9 0 5.7% 2895.38 1.32
TabM (T+E) 1529 0.439 12.8 8.5 0 4.3% 4228.53 1.19
CatBoost (T) $15$ 0.406 13.3 6.8 0 5.9% 2895.38 0.07
xRFM (T) 1504 0.406 13.8 9.7 0 5.7% 365.57 0.09
LightGBM (T) 1498 0.354 13.9 10.2 0 6.8% 686.46 0.74
XGBoost (T+E) 1470 0.328 15.2 13.9 0 6.7% 848.99 2.38
XGBoost (T) 1470 0.325 15.2 13.3 0 6.7% 848.99 0.47
ModernNCA (D) 1457 0.323 15.6 11.0 0 8.1% 16.07 0.29
TabM (T) 1453 0.311 16.0 12.6 0 5.2% 4228.53 0.13
RealMLP (T) 1439 0.306 16.6 13.3 0 5.9% 6860.54 0.32
CatBoost (D) 1419 0.278 17.4 12.0 0 7.9% 8.35 0.09
ModernNCA (T) 1411 0.347 17.6 6.7 0 7.9% 3811.43 0.45
TabPFNv2 (T+E) 1377 0.410 19.1 2.7 4 4.9% 3805.62 10.41
TabM (D) 1361 0.232 20.1 16.5 0 6.9% 13.90 0.12
TabPFNv2 (T) 1314 0.281 22.1 6.9 0 6.4% 3805.62 0.26
TorchMLP (T+E) 1301 0.141 22.7 19.1 0 8.1% 4452.11 0.85
ExtraTrees (T+E) 1297 0.169 23.1 16.9 0 11.1% 161.73 0.78
RealMLP (D) 1274 0.085 24.0 21.6 0 8.2% 23.30 1.44
ExtraTrees (T) 1273 0.174 24.0 16.7 0 11.4% 161.73 0.12
TabPFNv2 (D) 1264 0.262 24.3 7.7 0 7.6% 2.78 0.32
TorchMLP (T) 1245 0.124 25.2 20.8 0 8.7% 4452.11 0.09
xRFM (D) 1236 0.149 25.8 8.6 1 11.7% 1.65 0.08
LightGBM (D) 1231 0.069 25.9 24.5 0 9.7% 2.03 0.30
XGBoost (D) 1195 0.104 27.4 24.5 0 10.4% 2.15 0.18
RandomForest (T+E) 1193 0.056 27.5 25.5 0 12.1% 526.17 0.77
RandomForest (T) 1142 0.046 29.9 27.5 0 12.8% 526.17 0.12
EBM (T+E) 1133 0.122 30.0 16.4 0 14.7% 2124.78 0.12
ExtraTrees (D) 1117 0.056 30.6 27.7 0 13.0% 0.42 0.06
FastaiMLP (T+E) 1086 0.005 31.8 30.6 0 13.1% 527.21 2.83
EBM (T) 1076 0.131 32.2 9.4 1 15.3% 2124.78 0.01
TorchMLP (D) 1075 0.015 32.4 30.3 0 12.9% 20.50 0.08
FastaiMLP (T) 1046 0.000 33.3 32.3 0 13.6% 527.21 0.31
EBM (D) 1009 0.071 34.7 30.3 0 16.0% 7.25 0.04
TabICL (D) 1006 0.000 34.8 34.2 0 14.2% 0.63 0.06
RandomForest (D) 1000 0.000 34.8 34.2 0 14.2% 0.63 0.06
FastaiMLP (D) 916 0.000 37.2 36.5 0 17.8% 3.08 0.29
KNN (T+E) 533 0.000 43.8 43.6 0 37.3% 2.25 0.15
Linear (T+E) 485 0.000 44.3 44.2 0 35.5% 46.50 0.14
KNN (T) 434 0.000 44.9 44.7 0 38.0% 2.25 0.03
Linear (T) 426 0.000 44.9 44.8 0 35.7% 46.50 0.04
Linear (D) 312 0.000 46.0 46.0 0 38.1% 1.16 0.08
KNN (D) 263 0.000 46.5 46.2 0 41.6% 0.04 0.02
Table B.8: Full regression results on TabArena-Lite.
Model Elo ($\uparrow$) Norm. Avg. Harm. #wins ($\uparrow$) Improva- Train time Predict time
score ($\uparrow$) rank ($\downarrow$) mean bility ($\downarrow$) per 1K [s] per 1K [s]
rank ($\downarrow$)
AutoGluon 1.3 (4h) 1521 0.586 9.0 3.0 2 9.7% 4917.81 4.05
CatBoost (T) 1441 0.439 12.5 9.6 0 11.9% 3307.58 0.14
CatBoost (T+E) 1441 0.446 12.4 9.4 0 11.1% 3307.58 1.18
TabPFNv2 (T+E) 1419 0.518 13.5 4.2 1 13.8% 2584.13 12.37
LightGBM (T+E) 1418 0.393 13.2 7.5 0 13.6% 1280.01 4.08
LightGBM (T) 1405 0.381 13.9 6.9 0 13.6% 1280.01 1.05
TabM (T+E) 1387 0.405 14.8 9.7 0 15.7% 5568.31 1.78
XGBoost (T+E) 1377 0.310 15.2 13.3 0 14.0% 2029.77 4.11
TabM (T) 1373 0.401 15.5 9.1 0 16.0% 5568.31 0.37
XGBoost (T) 1373 0.317 15.6 13.1 0 13.7% 2029.77 1.04
RealMLP (T+E) 1359 0.337 16.2 5.3 1 16.8% 6866.35 10.40
TabPFNv2 (D) 1358 0.361 16.4 4.7 1 12.5% 5.48 0.35
xRFM (T+E) 1357 0.369 16.2 7.0 0 14.8% 515.01 1.68
TabPFNv2 (T) 1346 0.366 17.0 6.5 0 14.9% 2584.13 0.41
ModernNCA (T+E) 1345 0.302 16.9 10.4 0 15.3% 6684.65 9.59
ModernNCA (T) 1332 0.248 17.8 12.4 0 15.6% 6684.65 0.75
RealMLP (T) 1304 0.253 19.1 15.5 0 18.4% 6866.35 0.92
CatBoost (D) 1296 0.201 19.2 16.9 0 16.1% 43.10 0.25
TabICL (D) 1293 0.325 19.9 4.5 1 20.5% 11.51 1.95
TabM (D) 1293 0.284 19.9 9.4 0 19.7% 17.09 0.15
ExtraTrees (T+E) 1286 0.289 20.2 11.0 0 17.7% 728.32 2.44
xRFM (T) 1283 0.273 20.4 5.2 1 16.4% 515.01 0.20
ExtraTrees (T) 1265 0.262 21.2 11.2 0 17.3% 728.32 0.36
RandomForest (T+E) 1260 0.321 20.9 6.1 0 15.9% 729.17 1.83
TorchMLP (T+E) 1260 0.188 21.2 17.5 0 19.6% 3646.83 2.16
TorchMLP (T) 1218 0.156 23.9 17.8 0 21.6% 3646.83 0.19
RandomForest (T) 1190 0.241 24.9 12.1 0 17.6% 729.17 0.33
XGBoost (D) 1188 0.102 24.9 21.5 0 18.1% 4.93 0.59
TabDPT (D) 1186 0.259 25.2 4.7 1 23.3% 33.52 20.75
FastaiMLP (T+E) 1182 0.205 25.2 12.1 0 24.2% 2721.87 12.59
LightGBM (D) 1180 0.158 25.4 19.6 0 21.3% 5.12 0.44
EBM (T+E) 1168 0.166 26.0 17.1 0 24.6% 1471.12 0.27
EBM (T) 1144 0.160 27.4 19.0 0 24.9% 1471.12 0.03
xRFM (D) 1129 0.094 28.4 22.1 0 23.9% 2.22 0.20
ModernNCA (D) 1110 0.051 29.1 26.8 0 25.9% 17.24 0.57
FastaiMLP (T) 1100 0.104 29.6 20.3 0 25.8% 2721.87 1.08
RealMLP (D) 1049 0.071 32.0 24.5 0 26.3% 26.02 4.18
EBM (D) 1046 0.095 32.0 24.3 0 27.7% 6.16 0.08
RandomForest (D) 1000 0.000 33.9 32.0 0 33.3% 0.74 0.15
TorchMLP (D) 972 0.004 35.0 33.6 0 28.5% 14.37 0.36
FastaiMLP (D) 952 0.038 35.9 31.9 0 33.4% 8.37 0.66
ExtraTrees (D) 862 0.013 38.6 35.1 0 39.1% 0.76 0.15
Linear (T+E) 752 0.000 41.8 41.5 0 45.6% 170.51 0.20
Linear (T) 718 0.000 42.3 41.9 0 45.8% 170.51 0.13
KNN (T+E) 714 0.000 42.4 41.0 0 52.5% 2.99 0.17
Linear (D) 689 0.000 42.9 42.7 0 46.9% 3.89 0.16
KNN (D) 544 0.000 45.4 45.1 0 69.5% 0.33 0.05
KNN (T) 541 0.000 45.6 45.5 0 58.3% 2.99 0.06
Table B.9: Full multiclass results on TabArena-Lite.
Model Elo ($\uparrow$) Norm. Avg. Harm. #wins ($\uparrow$) Improva- Train time Predict time
score ($\uparrow$) rank ($\downarrow$) mean bility ($\downarrow$) per 1K [s] per 1K [s]
rank ($\downarrow$)
AutoGluon 1.3 (4h) 1779 0.673 5.2 3.2 1 3.6% 1734.20 7.06
RealMLP (T+E) 1721 0.660 6.6 5.4 0 3.1% 6860.54 7.68
ModernNCA (T+E) 1626 0.622 9.2 2.6 3 5.4% 3811.43 7.58
LightGBM (T+E) 1602 0.478 10.1 7.5 0 6.0% 686.46 5.48
TabDPT (D) 1575 0.515 10.9 3.8 2 3.9% 16.97 8.70
xRFM (T+E) 1563 0.492 11.5 5.3 1 5.1% 365.57 0.72
CatBoost (T+E) 1558 0.435 11.8 8.9 0 5.7% 2895.38 1.32
TabM (T+E) 1529 0.439 12.8 8.5 0 4.3% 4228.53 1.19
CatBoost (T) $15$ 0.406 13.3 6.8 0 5.9% 2895.38 0.07
xRFM (T) 1504 0.406 13.8 9.7 0 5.7% 365.57 0.09
LightGBM (T) 1498 0.354 13.9 10.2 0 6.8% 686.46 0.74
XGBoost (T+E) 1470 0.328 15.2 13.9 0 6.7% 848.99 2.38
XGBoost (T) 1470 0.325 15.2 13.3 0 6.7% 848.99 0.47
ModernNCA (D) 1457 0.323 15.6 11.0 0 8.1% 16.07 0.29
TabM (T) 1453 0.311 16.0 12.6 0 5.2% 4228.53 0.13
RealMLP (T) 1439 0.306 16.6 13.3 0 5.9% 6860.54 0.32
CatBoost (D) 1419 0.278 17.4 12.0 0 7.9% 8.35 0.09
ModernNCA (T) 1411 0.347 17.6 6.7 0 7.9% 3811.43 0.45
TabPFNv2 (T+E) 1377 0.410 19.1 2.7 4 4.9% 3805.62 10.41
TabM (D) 1361 0.232 20.1 16.5 0 6.9% 13.90 0.12
TabPFNv2 (T) 1314 0.281 22.1 6.9 0 6.4% 3805.62 0.26
TorchMLP (T+E) 1301 0.141 22.7 19.1 0 8.1% 4452.11 0.85
ExtraTrees (T+E) 1297 0.169 23.1 16.9 0 11.1% 161.73 0.78
RealMLP (D) 1274 0.085 24.0 21.6 0 8.2% 23.30 1.44
ExtraTrees (T) 1273 0.174 24.0 16.7 0 11.4% 161.73 0.12
TabPFNv2 (D) 1264 0.262 24.3 7.7 0 7.6% 2.78 0.32
TorchMLP (T) 1245 0.124 25.2 20.8 0 8.7% 4452.11 0.09
xRFM (D) 1236 0.149 25.8 8.6 1 11.7% 1.65 0.08
LightGBM (D) 1231 0.069 25.9 24.5 0 9.7% 2.03 0.30
XGBoost (D) 1195 0.104 27.4 24.5 0 10.4% 2.15 0.18
RandomForest (T+E) 1193 0.056 27.5 25.5 0 12.1% 526.17 0.77
RandomForest (T) 1142 0.046 29.9 27.5 0 12.8% 526.17 0.12
EBM (T+E) 1133 0.122 30.0 16.4 0 14.7% 2124.78 0.12
ExtraTrees (D) 1117 0.056 30.6 27.7 0 13.0% 0.42 0.06
FastaiMLP (T+E) 1086 0.005 31.8 30.6 0 13.1% 527.21 2.83
EBM (T) 1076 0.131 32.2 9.4 1 15.3% 2124.78 0.01
TorchMLP (D) 1075 0.015 32.4 30.3 0 12.9% 20.50 0.08
FastaiMLP (T) 1046 0.000 33.3 32.3 0 13.6% 527.21 0.31
EBM (D) 1009 0.071 34.7 30.3 0 16.0% 7.25 0.04
TabICL (D) 1006 0.000 34.8 34.2 0 14.2% 0.63 0.06
RandomForest (D) 1000 0.000 34.8 34.2 0 14.2% 0.63 0.06
FastaiMLP (D) 916 0.000 37.2 36.5 0 17.8% 3.08 0.29
KNN (T+E) 533 0.000 43.8 43.6 0 37.3% 2.25 0.15
Linear (T+E) 485 0.000 44.3 44.2 0 35.5% 46.50 0.14
KNN (T) 434 0.000 44.9 44.7 0 38.0% 2.25 0.03
Linear (T) 426 0.000 44.9 44.8 0 35.7% 46.50 0.04
Linear (D) 312 0.000 46.0 46.0 0 38.1% 1.16 0.08
KNN (D) 263 0.000 46.5 46.2 0 41.6% 0.04 0.02
Table B.10: Full binary classification results on TabArena.
@@
Figure B.1: Runtime comparison as a function of the number of training samples on the covertype dataset. Here, L1 kernel refers to the $L_{p}^{p}$ kernel.